A skateboarder rolls horizontally off the top of a staircase and lands at the bottom. The staircase has a horizontal length of 13 m, with a vertical drop of 3 m. We can ignore air resistance. What is the skater's initial horizontal velocity?



1
Expert's answer
2021-10-24T18:24:33-0400

The initial speed of the skater's

"v_0=\sqrt{2gh}=\sqrt{2*9.8*3}=19.6\:\rm m/s"